OpenStudy (anonymous):

Consider the quadratic relation y = x2 - 2x - 35. 1. Use the three different algebraic methods to rewrite the relation in vertex form.

OpenStudy (anonymous):

I don't know which three method but I will do what I know 1) By completing a square x^2-2x+1=35+1 (x-1)^2=36 (x-1)^2-36 2) Using -b/2a 2/2(1)=1 plug in 1 for x we get -36 for y (x-h)^2+k (x-1)^2-36

OpenStudy (anonymous):

THANKS!!!!!